OpenStudy (anonymous):

for what value of k the equation below has no solution?2x + 3 = x - 2kx - 5

hero (hero):

The equation will have no solution if x - 2kx = 2x solving for k: 2kx = x - 2x 2kx = -x 2kx + x = 0 x(2k + 1) = 0 2k + 1 = 0 2k = -1 k = -1/2 Therefore, if k = -1/2, then the equation 2x + 3 = x - 2kx - 5 will have no solution.